A potential field is given by V = 3x^2 y - yz. Which of the following is not true?

(a) At point (1,0,-1) both V and E vanish.

(b) x^2 y = 1 is an equipotential line on the xy-plane.

(c) The equipotential surface V = -8 passes through point P(2, -1, 4).

(d) The electric field at P (2,-1,4)is a 12ax - 8ay - az

(e) none of the above


1
Expert's answer
2021-07-07T08:41:09-0400

Gives

"V=3x^2y-yz"

"V_{(1,0,1)}=0"

"E=-\nabla V"



"E_{1,0,1}=-(6(0)a_x+(3\times1+1)a_y-0a_z)"

"E=-4a_y"

Option (a) wrong(incorrect)

"P=x^2y-1"

Equpotential line equation


"\nabla.P=\nabla.(x^2y-1)=2xya_x+x^2a_y\rightarrow(2)"


Equation (2) is lies xy-plane

Option (b) is correct option


Electric field

"V=3x^2y-yz"



"V_{(2,-1,4)}=3\times4(-1)-(-1)\times4=-8V"

Option (c) is correct



"E=-\nabla V=-(6xya_x+(3x^2-z)a_y-ya_z)"

"E=12a_x-8a_y-a_z"

Option (d) correct

"E=12ax-8ay-az"

"\hat{V}=\frac{\nabla V}{ |\nabla V|}"

"\hat{V}=\frac{6axy a_x+(3x^2-z)a_y+a_z(-y)}{14.5}"

"\hat{V}_{(2,-1,4)}=\frac{-12}{14.4}a_x+\frac{8}{14.4}a_y+\frac{1}{14.4}a_z"

"\hat{V}=-0.83a_x+0.55a_y+0.07a_z"

Option (e) is correct